Can you help me find the equation for the line?
(2,-5) and (0,4)
Could you explain it too?
Thank you

Answer by scott8148(6628)

first, find the slope
___ the slope is the difference in the y-coordinates divided by the difference in the x-coordinates
___ it is usually designated by the letter "m"

m = (-5 - 4) / (2 - 0) = -9/2 ___ remember to subtract the x and y coordinates in the same order

the point-slope form of a line is ___ y - [y-coordinate] = (slope) (x - [x-coordinate])
___ pick one of the points and use the coordinates ___ y - 4 = -9/2 (x - 0) ___ y = (-9/2)x